OpenStudy (anonymous):

divide this decimal 82.8 / 0.45

OpenStudy (anonymous):

If it is just basic division it's 184

OpenStudy (anonymous):

82.8/0.45 = 184

OpenStudy (anonymous):

\[\frac{ 82.80 }{ .45 }=\frac{ 8280 }{45 }\] now you can simplyfy

OpenStudy (anonymous):

mhmmm ok i see